# references.bib

### v3

@@ Abstract.
In this paper, we prove the integrality conjecture for quotient stacks given by a symmetric representation of a reductive group. This result will be crucial to generalise cohomological Donaldson–Thomas theory of $2$ and $3$-Calabi–Yau categories to 00 and $(-1)$-shifted symplectic stacks respectively.
@@ 1. Introduction
Given a connected reductive group $G$ and a representation $V$ of $G$, the quotient stack $\mathfrak{X}=V/G$ and the GIT quotient $V/\!\!/G$ are objects of interest in geometry and representation theory. In the context of quiver representations, the geometry and topology of the GIT quotient $V/\!\!/G$ has been studied deeply by Reineke, see [reineke2008moduli] for a survey. A particularly nice case is when $V$ is a symmetric representation of $G$, in the sense that the weights of $V$ counted with multiplicities are invariant under the involution $\alpha\mapsto-\alpha$.
@@
In [vspenko2017non, vspenko2021semi], the authors study semiorthogonal decompositions of the coherent derived category $\mathcal{D}(V/G)$ of $V/G$ in order to obtain noncommutative resolutions of $V/\!\!/G$. The tools are the window categories defined and studied by Halpern-Leistner [halpern2015derived]. Their formalism deals with quasi-symmetric representations $V$ of a reductive group $G$. Quasi-symmetricity is the vanishing property of the sum of weights of $V$ on any given line through the origin of the character lattice and is weaker than the symmetricity.
@@
In this paper, we are interested in the interplay between the topology of $V/G$ and that of $V/\!\!/G$. More precisely, we give a decomposition of the (infinite dimensional) cohomology vector space $\mathrm{H}^{*}(V/G)$ of the quotient stack $V/G$ in terms of a finite number of finite-dimensional cohomologically graded vector spaces $\mathcal{P}_{\lambda}$ indexed by cocharacters $\lambda\in X_{*}(T)$ of a maximal torus $T$ of $G$, using parabolic induction (Theorem 1.1), when $V$ is symmetric. Such a result is often called cohomological integrality. The refined Donaldson–Thomas invariant of $V/G$ are then defined as the Betti numbers of the vector spaces $\mathcal{P}_{\lambda}$. In this paper, we initiate the study of these enumerative invariants, which encode representation-theoretic information about $V/G$.
@@
In geometry, stacks of the form $V/G$ and their good moduli space $V/\!\!/G$ give, by Luna étale slice theorems [alper2020luna], local models for smooth stacks $\mathfrak{X}$ with a good moduli space $\mathcal{X}$. Therefore, this work is the ground for the study of the geometry and topology of morphisms $\mathfrak{X}\rightarrow\mathcal{X}$ from a symmetric smooth stack to its good moduli space.
When the group $G$ is a product of general linear groups and $V$ is a symmetric representation of $G$ given by the representation space of a symmetric quiver $Q$ (a quiver having as many arrows in both possible directions between any two vertices), the cohomological integrality isomorphism has been proven by Efimov in [efimov2012cohomological], answering a conjecture of Kontsevich and Soibelman [kontsevich2011cohomological, Conjecture 1]. It is formulated in terms of the cohomological Hall algebra of $Q$ [kontsevich2011cohomological]. The cohomological integrality of Efimov is the building block of all cohomological integrality results for categories of homological dimension one, and for $2$ and $3$-Calabi–Yau categories [meinhardt2019donaldson, davison2020cohomological, davison2022bps, davison2023bps]. These integrality theorems are one of the most sought-after theorems in the theory of cohomological Hall algebras, as they provide strong structural results and deep connections between the topology of the stack and the moduli space. The integrality also gives cohomologically refined enumerative invariants whose study is of algebro-geometric interest.
In particular, the cohomological integrality isomorphisms for the $2$-Calabi–Yau categories of semistable Higgs bundles on a smooth projective curve and twisted local systems is at the heart of nonabelian Hodge isomorphisms for stacks [davison2023nonabelian, davison2022bps, hennecart2023nonabelian]. Nonabelian Hodge isomorphisms for stacks are crucial to study the $\mathrm{P}=\mathrm{W}$ conjecture for $\mathrm{GL}_{n}$ for the singular moduli spaces.
A $2$-Calabi–Yau category $\mathscr{C}$ gives rise to a 00-shifted symplectic stack while a $3$-Calabi–Yau category $\mathscr{C}$ provides us with a $(-1)$-shifted symplectic stack [pantev2013shifted]. In the $2$-Calabi–Yau case, fully settled in [davison2022bps, davison2023bps], the cohomological integrality gives a decomposition of the Borel Moore homology $\mathrm{H}^{\mathrm{BM}}_{*}(\mathfrak{M}_{\mathscr{C}})$ of (a substack of) the stack of objects in the category $\mathscr{C}$ in terms of the intersection cohomology $\mathrm{IH}(\mathcal{M}_{\mathscr{C}})$ of the good moduli space $\mathcal{M}_{\mathscr{C}}$ of $\mathfrak{M}_{\mathscr{C}}$. The study of cohomological integrality for $2$-Calabi–Yau categories has several applications, such as the study of the cohomology of Nakajima quiver varieties and cuspidal polynomials of quivers [davison2023bps] and the comparison between cohomological Hall algebras and Maulik–Okounkov Yangians [maulik2019quantum, botta2023okounkov, schiffmann2023cohomological].
In the $3$-Calabi–Yau case, the cohomological integrality is, in full generality, a conjecture recalled in [davison2023bps, Conjecture 1.9]. It is proven for quivers with potential in [davison2020cohomological]. It concerns the study of the critical cohomology of the moduli stack of objects in the category. The critical cohomology is the cohomology of a perverse sheaf, which is either the sheaf of vanishing cycles or a globalisation of it [ben2015darboux, brav2015symmetries, brav2019darboux, kiem2012categorification]. In the case of general $3$-Calabi–Yau categories, significant progress towards the integrality conjecture is made in [kinjo2024cohomological].
All known cohomological integrality isomorphisms are expressed in terms of cohomological Hall algebras, which is a convenient and powerful way to gather the data of parabolic inductions. If rather than a $2$-Calabi–Yau category, one considers a 00-shifted symplectic stack $\mathfrak{M}$, one expects to have similar cohomological integrality isomorphisms. However, the cohomological Hall algebra structure is not available and one has to deal directly with the parabolic induction morphisms. This is what we do in this paper in the situation of smooth symmetric stacks given by a symmetric representation of a reductive group. In particular, the very definition of the cohomological integrality morphism (see Theorem 1.1) had to be given.
@@
There are motivic versions of integrality results, studied for example in [bu2024motivic] following [kontsevich2008stability] and motivic Donaldson–Thomas theory [kontsevich2010motivic, meinhardt2017introduction, bussi2019motivic].
@@ 1.1. Main results
In this paper, we work over an algebraically closed field of characteristic zero.
1.1.1. Induction
Let $G$ be a connected reductive group and $V$ a representation of $G$. We let $T\subseteq G$ be a maximal torus. For any cocharacter $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$, we let $G^{\lambda}\subseteq G$ be the corresponding Levi subgroup of $G$ [digne2020representations, Proposition 1.22] and $V^{\lambda}$ the representation of $G^{\lambda}$ given by the $\lambda$-fixed subspace of $V$. If $\lambda=0$ is the trivial cocharacter, we write $V=V^{0}$ and $G=G^{0}$.
@@ 1.1.1. Induction
where $d_{\lambda}=\dim V^{\lambda}-\dim G^{\lambda}$2.2). If $V$ is symmetric (i.e. $V$ and $V^{*}$ have the same sets of weights, Definition 2.3), then the induction map preserves cohomological degrees (Lemma 2.5).
@@
$\lambda\preceq\mu\stackrel{{\scriptstyle\text{def.}}}{{\iff}}\left\{\begin{aligned} &V^{\lambda}\subseteq V^{\mu}\\ &\mathfrak{g}^{\lambda}\subseteq\mathfrak{g}^{\mu}.\end{aligned}\right.$
We let $\mathscr{P}_{V}\coloneqq X_{*}(T)/\sim$ where $\sim$ is the equivalence relation given by $\lambda\sim\mu\iff(\lambda\preceq\mu\text{ and }\mu\preceq\lambda)$. In other words, two cocharacters are equivalent if their fixed-point sets inside $V$ and $\mathfrak{g}$ both coincide. The set $\mathscr{P}_{V}$ is finite. If $V$ is symmetric, the induction map $\mathrm{Ind}_{\lambda}$ only depends on the class $\overline{\lambda}\in\mathscr{P}_{V}$ up to a sign (Lemma 2.14).
@@ 1.1.2. Cohomological integrality
Let $G$ be a reductive group with maximal torus $T$ and $V$ be a finite dimensional symmetric representation of $G$ (Definition 2.3). For any $\overline{\lambda}\in\mathscr{P}_{V}$, we let $\lambda\in X_{*}(T)$ be an arbitrary lift. For $\widetilde{\lambda}\in\mathscr{P}_{V}/W$, we let $\lambda\in X_{*}(T)$ be an arbitrary lift.
@@
We let $G_{\lambda}\subseteq G$ be the intersection of the kernel of the action map $G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda})$ with the center of $G^{\lambda}$ and $\mathfrak{g}_{\lambda}\coloneqq\mathrm{Lie}(G_{\lambda})$. Note that the group $G_{\lambda}$ is not necessarily connected, but this can essentially be ignored (Lemma 4.19).
For $\alpha\in X^{*}(T)$, we let $V_{\alpha}\coloneqq\{v\in V\mid\forall t\in T,t\cdot v=\alpha(t)\}$ be the $\alpha$-weight space of $V$. We define $\mathfrak{g}_{\alpha}$ similarly.
@@
$k_{\lambda}\coloneqq\frac{\prod_{\alpha\in X^{*}(T),\langle\lambda,\alpha\rangle>0}\alpha^{\dim V_{\alpha}}}{\prod_{\alpha\in X^{*}(T),\langle\lambda,\alpha\rangle>0}\alpha^{\dim\mathfrak{g}_{\alpha}}}.$
We may call $k_{\lambda}$ the induction kernel. We define $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\{\pm 1\}$ to be the sign such that for any $w\in W_{\lambda}$, $w(k_{\lambda})=\varepsilon_{V,\lambda}(w)k_{\lambda}$ (Proposition 4.8).
@@
There exists cohomologically graded vector spaces $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ with a $W_{\lambda}$-action, $\lambda\in X_{*}(T)$, such that the map
@@
induced by the induction morphisms (1.1), is a graded isomorphism, where $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\cong\mathrm{Sym}(\mathfrak{g}_{\lambda}^{*})$, $\mathfrak{g}_{\lambda}^{*}$ sits in degree $2$, $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\{\pm 1\}$ is a character and $(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}))^{\varepsilon_{V,\lambda}}$ denotes the $\varepsilon_{V,\lambda}$-isotypic component for the natural $W_{\lambda}$-action. Moreover, the vector spaces $\mathcal{P}_{\lambda}$ are finite-dimensional and the inclusion $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ factors through the inclusion $\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))\rightarrow\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$.
This type of theorems is sometimes referred to as cohomological integrality. We can reconstruct the infinite dimensional vector space $\mathrm{H}^{*}(V/G)$ using a finite number of finite-dimensional vector spaces $\mathcal{P}_{\lambda}$, $\widetilde{\lambda}\in\mathscr{P}_{V}/W$.
@@
Let $Q=(Q_{0},Q_{1})$ be a symmetric quiver with set of vertices $Q_{0}$ and set of arrows $Q_{1}$. Let $\mathbf{d}\in\mathbf{N}^{Q_{0}}$ be a dimension vector. If $V_{\mathbf{d}}$ is the representation space of $\mathbf{d}$-dimensional representations of $Q$ which is acted on by $\mathrm{GL}_{\mathbf{d}}\coloneqq\prod_{i\in Q_{0}}\mathrm{GL}_{\mathbf{d}_{i}}$, then the isomorphism given by Theorem 1.1 essentially recovers [efimov2012cohomological, Theorem 1.1]. In loc.cit, it was convenient for the author to twist the cohomological Hall algebra multiplication to make it supercommutative, as in [kontsevich2011cohomological]. This twist does not change the images of the induction/multiplication morphisms which means that it is possible to have cohomological integrality isomorphisms even without this twist, although we cannot identify the untwisted CoHA with a supercommutative algebra anymore. We can only identify the underlying vector spaces.
@@
We define the cohomologically refined Donaldson–Thomas invariants of $V/G$ as the dimensions $p_{\lambda,i}\coloneqq\dim\mathcal{P}_{\lambda}^{i}$ of the cohomological degree $i$ piece of the vector space $\mathcal{P}_{\lambda}$ for $\lambda\in X_{*}(T)$. We may call the Euler characteristics $p_{\lambda}\coloneqq\sum_{i\in\mathbf{Z}}(-1)^{i}p_{\lambda,i}$ of $\mathcal{P}_{\lambda}$ the Donaldson–Thomas invariants of $V/G$.

### v6

# Cohomological integrality for weakly symmetric representations of reductive groups

@@ Abstract.
In this paper, we prove the integrality conjecture for quotient stacks arising from weakly symmetric representations of reductive groups. Our main result is a decomposition of the cohomology of the stack into finite-dimensional components indexed by some equivalence classes of cocharacters of a maximal torus. This decomposition enables the definition of new enumerative invariants associated with the stack, which we begin to explore.
1991 Mathematics Subject Classification:
14L30, 14L24, 13A50
@@ 1. Introduction
1.1. Invariant Theory
Given a connected reductive group $G$ and a finite-dimensional representation $V$ of $G$, the quotient stack $\mathfrak{X}=V/G$ and the Geometric Invariant Theory (GIT) quotient $\mathcal{X}=V/\!\!/G$ are fundamental objects in invariant theory [MFK94, DC70, PV94], with significant applications to geometry and representation theory, e.g. [Nak94, Rei08, MO19] for connections with Lie theory.
@@ 1.1. Invariant Theory
Recent research has continued to explore the geometry of the quotient $V/G$. For instance, in [ŠV17, ŠV21], the authors study semiorthogonal decompositions of the coherent derived category $\mathcal{D}(V/G)$ to obtain noncommutative resolutions of V//GV/\!\!/G [Van04, Lun10, KL15], using the framework of window categories developed by Halpern-Leistner [Hal15]. Their approach focuses on quasi-symmetric representations of reductive groups, which are the representations for which the sum of weights on any line through the origin in the character lattice is zero.
@@
In this paper, we consider the case where $V$ is a weakly symmetric representation of $G$ (Definition 2.3). This encompasses the case of self-dual representations. We prove a decomposition of the (infinite-dimensional) cohomology vector space $\mathrm{H}^{*}(V/G)$ into a finite collection of finite-dimensional, cohomologically graded vector spaces $\mathcal{P}_{\lambda}$ indexed by cocharacters $\lambda\in X_{*}(T)$ of a maximal torus $T$ of $G$ (Theorem 1.1). This decomposition, which employs parabolic induction (§2), is known as cohomological integrality and can be seen as a cohomological version of a generalised Springer decomposition [Lus84] in this context. We then define the refined Donaldson–Thomas invariants of $V/G$ as the Betti numbers of these vector spaces $\mathcal{P}_{\lambda}$ (Definition 1.4) following [MR19, Mei15, DM20] and initiate their study (Proposition 1.5). These invariants are expected to encode rich representation-theoretic information about the action of $G$ on $V$ and topological information about the GIT quotient V//GV/\!\!/G. In particular, we expect that the vector space $\mathcal{P}_{\lambda}$ computes the intersection cohomology of some GIT quotient, thereby providing an efficient algorithmic and inductive way to determine the intersection Betti numbers of a large class of GIT quotients. We study this question in [Hen24] and in forthcoming work.
@@
Stacks of the form $V/G$ and their moduli spaces V//GV/\!\!/G serve as local models for smooth stacks $\mathfrak{X}$ with good moduli spaces $\mathcal{X}$ by the Luna étale slice theorem [AHR20]. Therefore, our results provide a foundation for exploring the geometry and topology of morphisms $\mathfrak{X}\rightarrow\mathcal{X}$ from a weakly symmetric smooth stack to its good moduli space [Alp13]. While this paper develops the core algebraic results, the applications to the geometry of a class of significant stacks in algebraic geometry and representation theory is investigated in [Hen24].
1.2. Cohomological Integrality
Cohomological integrality results, originally developed by Kontsevich and Soibelman in the context of cohomological Hall algebras and wall-crossing for moduli spaces [KS08, KS11], involve factorizations of generating series into plethystic exponentials. For symmetric quivers without potential, Efimov [Efi12] proved a cohomological integrality isomorphism, answering a conjecture by Kontsevich and Soibelman [KS11], which identifies the cohomological Hall algebra (CoHA) to a supercommutative algebra over a graded vector space with finite-dimensional graded components. Efimov’s isomorphism can be seen as a categorification of integrality formulas, in the sense that an equality between two generating series is turned into an isomorphism between two graded vector spaces whose graded characters recover the series.
Efimov’s integrality theorem forms the basis of all subsequent cohomological integrality results for CoHAs [MR19, DM20, KV22, DHM22, DHM23]. These results provide deep structural insights and link the topology of the stack to the good moduli space. In this context, cohomological integrality is often expressed through Poincaré-Birkhoff-Witt (PBW) theorems, analogous to those in the theory of Lie algebras, and leads to refined cohomological invariants of interest in algebraic geometry, forming the basis of cohomological Donaldson–Thomas theory [Sze14, DM20, Dav23].
In contrast to the traditional approach involving cohomological Hall algebras, our work deals directly with parabolic induction morphisms (§2) for more general smooth weakly symmetric stacks arising from weakly symmetric representations of reductive groups. We define the cohomological integrality morphism explicitly in Theorem 1.1. It does not involve the symmetric power of a vector space as in the quiver case, but some combinatorial replacement of it for any Weyl group, defined from the lattice of cocharacter of a maximal torus.
This paper can be viewed as a generalization of Efimov’s integrality isomorphism [Efi12] to all weakly symmetric stacks of the form $V/G$, not just those arising from quivers. It lays the groundwork for extending cohomological Donaldson–Thomas theory to more general stacks and moduli spaces beyond those parameterizing objects in an Abelian category. The cohomological integrality we present here is a purely algebraic result, with a purely algebraic proof. We explain in [Hen24] the geometric implications of this result.
@@ 1.2. Cohomological Integrality
There are also motivic versions of integrality results, such as those studied in [Bu24] following [KS08] and in the context of motivic Donaldson–Thomas theory [KS10, Mei17, BJM19], which are different from ours.
@@ 1.3. Main results
In this paper, we work over an algebraically closed field of characteristic zero. Moreover, cohomology is considered with $\mathbf{Q}$-coefficients.
1.3.1. Parabolic induction
We refer to §2.1 for the precise definitions. Let $G$ be a connected reductive group and $V$ a representation of $G$. We let $T\subseteq G$ be a maximal torus. For any cocharacter $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$, we let $G^{\lambda}\subseteq G$ be the corresponding Levi subgroup of $G$ [DM20a, Proposition 1.22] and $V^{\lambda}$ the representation of $G^{\lambda}$ given by the $\lambda$-fixed subspace of $V$. If $\lambda=0$ is the trivial cocharacter, we write $V=V^{0}$ and $G=G^{0}$.
@@ 1.3.1. Parabolic induction
where $d_{\lambda}=\dim V^{\lambda}-\dim G^{\lambda}$2.2). If $V$ is symmetric (i.e. $V$ and $V^{*}$ have the same sets of weights, Definition 2.3), then the induction map preserves shifted cohomological degrees (Lemma 2.7), that is $\mathrm{Ind}_{\lambda}(\mathrm{H}^{i+d_{\lambda}}(V^{\lambda}/G^{\lambda}))\subseteq\mathrm{H}^{i+d_{0}}(V/G)$ for any $i\in\mathbf{Z}$ (this is the reason for the notation $\mathrm{H}^{*+d_{\lambda}}$ in (1.1), which emphasizes that the natural grading is the cohomological grading shifted by $d_{\lambda}$).
@@
(1.2) $\lambda\preceq\mu\stackrel{{\scriptstyle\text{def.}}}{{\iff}}\left\{\begin{aligned} &V^{\lambda}\subseteq V^{\mu}\\ &\mathfrak{g}^{\lambda}\subseteq\mathfrak{g}^{\mu}\,.\end{aligned}\right$.
We let $\mathscr{P}_{V}\coloneqq X_{*}(T)/\sim$, where $\sim$ is the equivalence relation given by $\lambda\sim\mu\iff(\lambda\preceq\mu\text{ and }\mu\preceq\lambda)$. In other words, two cocharacters are equivalent if their fixed-point sets inside $V$ and $\mathfrak{g}$ (or equivalently inside $V$ and $G$) both coincide. The set $\mathscr{P}_{V}$ is finite. If $V$ is symmetric, the induction map $\mathrm{Ind}_{\lambda}$ only depends on the class $\overline{\lambda}\in\mathscr{P}_{V}$ up to a non-zero scalar (Lemma 2.16).
@@ 1.3.2. Cohomological integrality
Let $G$ be a reductive group with maximal torus $T$ and $V$ be a finite dimensional weakly symmetric representation of $G$ (Definition 2.3). For any $\overline{\lambda}\in\mathscr{P}_{V}$, we let $\lambda\in X_{*}(T)$ be an arbitrary lift. For $\widetilde{\lambda}\in\mathscr{P}_{V}/W$, we let $\lambda\in X_{*}(T)$ be an arbitrary lift.
@@
We let $G_{\lambda}\subseteq G$ be the intersection of the kernel of the action map $G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda})$ with the center of $G^{\lambda}$ and $\mathfrak{g}_{\lambda}\coloneqq\mathrm{Lie}(G_{\lambda})$. Note that the group $G_{\lambda}$ is not necessarily connected, but this can essentially be ignored (Lemma 4.1) because we consider cohomology with $\mathbf{Q}$-coefficients.
For $\alpha\in X^{*}(T)$, we let $V_{\alpha}\coloneqq\{v\in V\mid\forall t\in T,t\cdot v=\alpha(t)v\}$ be the $\alpha$-weight space of $V$. We define $\mathfrak{g}_{\alpha}$ similarly.
@@
$k_{\lambda}\coloneqq\frac{\prod_{\alpha\in X^{*}(T),\langle\lambda,\alpha\rangle<0}\alpha^{\dim V_{\alpha}}}{\prod_{\alpha\in X^{*}(T),\langle\lambda,\alpha\rangle<0}\alpha^{\dim\mathfrak{g}_{\alpha}}}\,.$
We may call $k_{\lambda}$ the induction kernel, see the shuffle-like formula in Proposition 2.11. We define $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\mathbf{C}^{*}$ to be the character such that for any $w\in W_{\lambda}$, $w(k_{\lambda})=\varepsilon_{V,\lambda}(w)^{-1}k_{\lambda}$ (Proposition 4.6).
@@
There exist $W_{\lambda}$-stable cohomologically graded vector spaces $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, $\lambda\in X_{*}(T)$, such that the map
@@
induced by the induction morphisms (1.1) is a graded isomorphism, where $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\cong\mathrm{Sym}(\mathfrak{g}_{\lambda}^{*})$, $\mathfrak{g}_{\lambda}^{*}$ sits in degree $2$ and $(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}))^{\varepsilon_{V,\lambda}}$ denotes the $\varepsilon_{V,\lambda}$-isotypic component for the natural $W_{\lambda}$-action. Moreover, the vector spaces $\mathcal{P}_{\lambda}$ are finite-dimensional and the inclusion $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ factors through the inclusion $\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))\rightarrow\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$. In addition, the vector spaces $\mathcal{P}_{\mu}$ for $\mu\preceq\lambda$ satisfy the cohomological integrality isomorphism for $(V^{\lambda},G^{\lambda})$.
This type of theorems is sometimes referred to as cohomological integrality. We can reconstruct the infinite dimensional vector space $\mathrm{H}^{*}(V/G)$ using a finite number of finite-dimensional vector spaces $\mathcal{P}_{\lambda}$, $\widetilde{\lambda}\in\mathscr{P}_{V}/W$, allowing us to extract fundamental enumerative invariants associated to the stack $V/G$ (Definition 1.4).
For a cocharacter $\lambda\in X_{*}(T)$, there is a natural $W$-equivariant map $\mathscr{P}_{V^{\lambda}}\rightarrow\mathscr{P}_{V}$, which can be defined as follows. Given a cocharacter $\mu\in X_{*}(T)$, there exists a cocharacter $\nu\in X_{*}(T)$ such that $V^{\lambda}\cap V^{\mu}=V^{\nu}$ and $G^{\lambda}\cap G^{\mu}=G^{\nu}$ (Lemma 2.8). We have $\nu\preceq\lambda$ and moreover, $\overline{\nu}=\overline{\mu}$ in $\mathscr{P}_{V^{\lambda}}$. The map sends $\overline{\mu}=\overline{\nu}\in\mathscr{P}_{V^{\lambda}}$ to $\overline{\nu}\in\mathscr{P}_{V}$. Then, the last sentence of Theorem 1.1 means that the vector spaces $\mathcal{P}_{\nu}$ satisfy the cohomological integrality isomorphism for $(V^{\lambda},G^{\lambda})$.
@@
Let $Q=(Q_{0},Q_{1})$ be a symmetric quiver with set of vertices $Q_{0}$ and set of arrows $Q_{1}$. Let $\mathbf{d}\in\mathbf{N}^{Q_{0}}$ be a dimension vector. If $V_{\mathbf{d}}$ is the representation space of $\mathbf{d}$-dimensional representations of $Q$ which is acted on by $\mathrm{GL}_{\mathbf{d}}\coloneqq\prod_{i\in Q_{0}}\mathrm{GL}_{\mathbf{d}_{i}}$, then the isomorphism given by Theorem 1.1 essentially recovers [Efi12, Theorem 1.1]. In loc.cit, it was convenient for the author to twist the cohomological Hall algebra multiplication to make it supercommutative, as in [KS11]. This twist does not change the images of the induction/multiplication morphisms which means that it is possible to have cohomological integrality isomorphisms even without this twist. This remark inspired the present work, in which the twist is replaced by the character $\varepsilon_{V,\lambda}$ of the group $W_{\lambda}$ (Proposition 4.6). When $V$ is symmetric (and not only weakly symmetric), the character $\varepsilon_{V,\lambda}$ takes values in $\{\pm 1\}$.
Remark 1.3.
In this paper, we consider cohomology with $\mathbf{Q}$-coefficients, which is crucial for the cohomological integrality map (Theorem 1.1) to be an isomorphism. It is crucial to work with $\mathbf{Q}$-coefficients from §2.5 until the end of the paper. However, one can define the induction morphisms (§2) for any ring of coefficients. It then becomes an interesting question to study the analogous question of cohomological integrality in positive characteristic, especially when the cohomology of the classifying space $\mathrm{H}^{*}(\mathrm{pt}/G)$ has torsion.
@@
We define the cohomologically refined Donaldson–Thomas invariants of $V/G$ as the dimensions $p_{\lambda,i}\coloneqq\dim\mathcal{P}_{\lambda}^{i}\subseteq\mathrm{H}^{i+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ of the cohomological degree $i$ piece of the vector space $\mathcal{P}_{\lambda}$ for $\lambda\in X_{*}(T)$. We may call the Euler characteristics $p_{\lambda}\coloneqq\sum_{i\in\mathbf{Z}}(-1)^{i}p_{\lambda,i}$ of $\mathcal{P}_{\lambda}$ the Donaldson–Thomas invariants of $V/G$.